OpenStudy (anonymous):

A particle moves according to the equation x = 11t2, where x is in meters and t is in seconds (a) Find the average velocity for the time interval from 2.20 s to 2.95 s. (b) Find the average velocity for the time interval from 2.20 s to 2.40 s.

OpenStudy (anonymous):

v(average)=(x2-x1)/(t2-t1) since x2=11*2.95*2.95 and x1=11*2.20*2.20 where t1=2.95, t1=2.20 put those values into equation and u ll get the answer

OpenStudy (anonymous):

Thanks :)